OpenStudy (anonymous):

write an equation of the line, in point-slope form, that passes through points (-15,7) and 5,-3) write an equation in slope-intercept form. What are the slope and y-intercept? -3x-10y=7

OpenStudy (shinalcantara):

Point Slope Form: \[(y - y _{1}) = m(x - x _{1})\] Slope Formula: \[m = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] You need to find first the slope using the slope formula Solve for slope

OpenStudy (anonymous):

this makes no sense.:/

OpenStudy (anonymous):

wait would the answer be y-7=-1/2(x+15) ??

OpenStudy (nikato):

correct!

OpenStudy (anonymous):

oh, thank you! for the second question the answer would be y=-3/10x-7/10 slope: -3/10 y intercept: -7/10 ????

OpenStudy (nikato):

correct again!

OpenStudy (anonymous):

alright, thank you both sooooo much!!:)

OpenStudy (nikato):

no problem

OpenStudy (shinalcantara):

if you're given two points of the equation, all you have to do is find first the slope of the line \[m = \frac{ -3 - 7 }{ 5 - (-15) }\] \[m = \frac{ -10 }{ 20 }\] \[m = \frac{ -1 }{ 2 }\] Since you already have the slope of the line, all you have to do is plug in any of the two points given (-15,7) or (5,-3) to the point slope form in this case lemme use (5,-3) plugging in you will have: \[(y - (-3)) = -\frac{ 1 }{ 2 }(x - 5)\] Since you are to have it in the point slope form then \[y + 3 = -\frac{ 1 }{ 2 }(x - 5)\]

OpenStudy (anonymous):

alright, thank you so much :)